# Implicit Hypergraph Neural Network

### v1

@@ IV-C Experimental Setup
We conducted all experiments on AMD EPYC 7763 64-Core Processor with 1.08 TB of memory and 8 NVIDIA A40 GPUs with CUDA version 12.1. Our code and experimental setup, including data construction, are available for peer review. For all the datasets, the train: test split is 30:70, and each model’s performance is averaged over 10 independent runs containing random data splits. For experiments on DBLP and Cora, all models are trained for 200 epochs while High-school and House-Bills are trained for 1000 epochs.
@@ V-A RQ1: Transductive Node Classification
Our first experiment is a simple transductive node classification on the five datasets described earlier. The results are summarized in Table III. Overall, the results demonstrate that IHNN outperforms all the state-of-the-art baselines with an overall rank of 1.2, while the closest baseline has the overall rank of 2.6. This result demonstrates that IHNN is able to incorporate both node and edge information better than the baselines.
Since our experiment involves diverse datasets from different application domains, there are several interesting observations we make. First, let us consider the co-authorship networks, Cora and DBLP. AllSet Transformer, which is a generalized message-pooling mechanism, outperforms HGNN by 9.45% on DBLP, while having a similar performance on Cora. However, IHNN outperforms AllSet by 7.80% on DBLP and 4.41% on Cora. Second, we notice that MLP outperforms CI on these two datasets but has worse performance for the 3 other datasets. This is because the co-authorship datasets have initial node features computed from the data while the initial node features were randomly initialized for the 3 other datasets. So, CI which uses the hypergraph structure to form node representations is not affected by input data noise like MLP in the 3 other datasets. Thirdly, we observe that IHNN significantly outperforms the nearest baseline on datasets where the number of hyperedges is much larger than the number of nodes (DBLP, High-school, House-bills). This demonstrates the ability of IHNN to capture long-range dependencies between nodes in hypergraphs. This empirical result is similar to the observation [14] previously made for standard graphs. Overall, IHNN outperforms the nearest baseline on 4 out of the 5 datasets.
@@ V-B RQ2: Ablation Study
The results are presented in Figure 3. We display the mean test accuracy for the 30:70 train: test split averaged across 10 independent runs for each component drop. First of all, we notice that abandoning the implicit form results in about a 15.99% drop in the accuracy of the overall framework, highlighting the need for an implicit model. Additionally, we highlight the importance of the cyclic form of our model and initial edge embeddings by randomly initializing edge features, which results in about a 32.90% drop in performance. Finally, the random initialization of the node features in IHNN leads to a 65.87% drop in performance. This shows the importance of the bias term (initial node features) to properly guide the solution of a fixed-point equilibrium equation involving implicitly defined vectors.
@@ V-D RQ4: Case Study: Uncovering Medical Code Relationships from EHR Data
To demonstrate the broader applicability of IHNN, we perform a case study on real-world health data. Hypergraphs have been used to encode the relationship between medical codes in Electronic Health Records (EHR) and also have been to improve the prediction of patient risks at an individual level [22, 2]. Our case study focuses on leveraging IHNN to estimate patient risks and then analyzing the inferred relationships between medical codes.
We begin by extracting inpatient medical information and ICD-9 billing codes for patients admitted between January 1, 2011, and June 30, 2011 from a real-world EHR dataset (R-EHR) collected from a large tertiary-care healthcare facility. Each patient visit is typically associated with multiple ICD-9 codes indicating multiple diagnoses during the hospital stay. ICD-9 codes are obtained from the patient’s discharge summary. Our risk prediction task for this case study is readmission prediction. We constructed binary labels indicating whether the patient was readmitted to the hospital within 30 days of discharge. The label statistics are provided in Table IV.
@@
TABLE IV: Readmission Counts Statistics for the case study.
In the hypergraph we constructed, each unique patient visit to the healthcare facility is considered a node, and ICD-9 codes serve as hyperedges. This resulted in a hypergraph consisting of 8435 nodes and 3375 hyperedges. We trained IHNN with node classification loss on the readmission label. We randomly split visits to 30:70 train: test ratio. Note that the label imbalance makes the ROC-AUC score a more suitable metric for evaluating this predictive performance. IHNN obtained a test ROC-AUC of 68.39 while the best-baseline HGNN obtained an ROC-AUC of 62.64.
Here, we are also interested in understanding if the relations between the hyper-edges (ICD-9 codes) inferred by IHNN correspond to domain knowledge. Firstly, we mapped the 3375 ICD-9 codes to 274 broad Clinical Classification Software (CCS) concepts. We then used the number of unique CCS concepts as the number of clusters to perform spectral clustering on the hyperedge embeddings and display the top 10 largest (in terms of number of elements) clusters with T-SNE visualization in Figure 5. In the figure, the purple cluster is distant from the rest. On manual inspection, we found that the cluster contains the ICD-9 codes referring to adverse effects & poisoning, cardiovascular diseases, respiratory, renal, neurological, gastrointestinal, urinary, and endocrine disorders. Detailed Descriptions of the meanings of these codes are given in Table V.
@@
TABLE V: ICD-9 Codes and Their Descriptions for the purple points.
Amongst these codes, we observed that eight of them are directly or indirectly related to heart attack. 411.89 is directly related, as ischemic heart disease [41] often leads to myocardial infarction. 429.9 is also directly linked, as it encompasses various heart conditions, including those leading to heart attacks. Indirectly related codes include 428.20 since heart failure can result from repeated heart attacks or prolonged ischemia; 398.90, as valve damage from rheumatic fever increases the risk of ischemic events [42]; 440.21 and 440.4 since atherosclerosis is a major cause of coronary artery disease and heart attacks [43]; 429.83, Takotsubo syndrome, which mimics a heart attack and is triggered by extreme stress [44]; and 996.61, as infections from cardiac implants can lead to complications affecting heart function [45]. These conditions either directly cause myocardial infarction or indicate underlying cardiovascular issues that increase heart attack risk.
@@
TABLE VI: ICD-9 Codes and Their Descriptions for the teal points.
Similarly, the cluster of points colored in teal primarily contains ICD-9 codes associated with mental health and neurological issues. A detailed description of the ICD-9 codes and their meanings is given in Table VI. 386.11 is directly related, as benign paroxysmal positional vertigo (BPPV) affects the vestibular system, leading to dizziness and balance issues [46]. 357.0 and 357.4 involve polyneuropathies [47], which affect the peripheral nervous system and can result in weakness, numbness, or pain. 784.69 is linked to speech disturbances, often indicative of neurological disorders. Mental health conditions are represented by 296.44 and 296.42, which describe different severities of bipolar I disorder, impacting mood and cognition. 295.40 covers schizophreniform disorder, a condition with symptoms resembling schizophrenia. Substance-related disorders, such as 305.43 (opioid abuse with other drug abuse) and 305.30 (cannabis abuse), highlight the intersection between mental health and addiction. 437.8, which includes cerebrovascular diseases, may contribute to cognitive impairment and neurological deficits. These conditions directly or indirectly affect brain function, mental well-being, and the nervous system.

### v2

@@ IV-C Experimental Setup
We conducted all experiments on AMD EPYC 7763 64-Core Processor with 1.08 TB of memory and 8 NVIDIA A40 GPUs with CUDA version 12.1. Our code and experimental setup, including data construction, are available for peer review.  [^1] https://github.com/Soothysay/IHNN. For all the datasets, the train: test split is 30:70, and each model’s performance is averaged over 10 independent runs containing random data splits. For experiments on DBLP and Cora, all models are trained for 200 epochs, while High-school and House-Bills are trained for 1000 epochs.
@@ V-A RQ1: Transductive Node Classification
Our first experiment is a simple transductive node classification on the five datasets described earlier. The results are summarized in Table II. Overall, the results demonstrate that IHNN outperforms all the state-of-the-art baselines with an overall rank of 1.2, while the closest baseline has the overall rank of 2.6. This result demonstrates that IHNN is able to incorporate both node and edge information better than the baselines.
Since our experiment involves diverse datasets from different application domains, there are several interesting observations we make. First, let us consider the co-authorship networks, Cora and DBLP. AllSet Transformer, which is a generalized message-pooling mechanism, outperforms HGNN by 9.45% on DBLP, while having a similar performance on Cora. However, IHNN outperforms AllSet by 7.80% on DBLP and 4.41% on Cora. Second, we notice that MLP outperforms CI on these two datasets but has worse performance for the 3 other datasets. This is because the co-authorship datasets have initial node features computed from the data while the initial node features were randomly initialized for the 3 other datasets. So, CI which uses the hypergraph structure to form node representations is not affected by input data noise like MLP in the 3 other datasets. Thirdly, we observe that IHNN significantly outperforms the nearest baseline on datasets where the number of hyperedges is much larger than the number of nodes (DBLP, High-school, House-bills). This demonstrates the ability of IHNN to capture long-range dependencies between nodes in hypergraphs. This empirical result is similar to the observation [gu2020implicit] previously made for standard graphs. Overall, IHNN outperforms the nearest baseline on 4 out of the 5 datasets.
@@ V-B RQ2: Ablation Study
The results are presented in Figure 3. We display the mean test accuracy for the 30:70 train: test split averaged across 10 independent runs for each component drop. First of all, we notice that abandoning the implicit form (w/o implicit) results in about a 15.99% drop in the accuracy of the overall framework, highlighting the need for an implicit model. Additionally, we highlight the importance of the cyclic form of our model and initial edge embeddings by randomly initializing edge features, which results in about a 32.90% drop in performance. Finally, using IHNN without the membership regularizer (w/o mem reg) leads to a 1.39% drop in performance. This shows that each individual component of our overall framework is important for the gain in performance.
@@ V-D RQ4: Case Study: Uncovering Medical Code Relationships from EHR Data
To demonstrate the broader applicability of IHNN, we perform a case study on real-world health data. Hypergraphs have been used to encode the relationship between medical codes in Electronic Health Records (EHR) and also have been to improve the prediction of patient risks at an individual level [xu2023hypergraph, choudhuri2025domain]. Our case study focuses on leveraging IHNN to estimate patient risks and then analyzing the inferred relationships between medical codes.
We begin by extracting inpatient medical information and ICD-9 billing codes for patients admitted between January 1, 2011, and June 30, 2011 from a real-world EHR dataset (R-EHR) collected from the University of Iowa Hospitals and Clinics (UIHC) [jang2022risk, choudhuri2023continually]. Each patient visit is typically associated with multiple ICD-9 codes indicating multiple diagnoses during the hospital stay. ICD-9 codes are obtained from the patient’s discharge summary. Our risk prediction task for this case study is readmission prediction. We constructed binary labels indicating whether the patient was readmitted to the hospital within 30 days of discharge. The label statistics are provided in Table III.
@@
TABLE III: Readmission Counts Statistics for the case study.
In the hypergraph we constructed, each unique patient visit to the healthcare facility is considered a node, and ICD-9 codes serve as hyperedges. This resulted in a hypergraph consisting of 8435 nodes and 3375 hyperedges. We trained IHNN with node classification loss on the readmission label. We randomly split visits to a 30:70 train: test ratio. Note that the label imbalance makes the ROC-AUC score a more suitable metric for evaluating this predictive performance [choudhuri2025domain]. IHNN obtained a test ROC-AUC of 68.39. Here, we are also interested in understanding if the relations between the hyper-edges (ICD-9 codes) inferred by IHNN correspond to domain knowledge. Firstly, we mapped the 3375 ICD-9 codes to 274 broad Clinical Classification Software (CCS) concepts. We then used the number of unique CCS concepts as the number of clusters to perform spectral clustering on the hyperedge embeddings and display the top 10 largest (in terms of number of elements) clusters with T-SNE visualization in Figure 5. In the figure, the purple cluster is distant from the rest. On manual inspection, we found that the cluster contains the ICD-9 codes referring to adverse effects & poisoning, cardiovascular diseases, respiratory, renal, neurological, gastrointestinal, urinary, and endocrine disorders. Detailed Descriptions of the meanings of these codes are given in Table IV.
@@
TABLE IV: ICD-9 Codes and Their Descriptions for the purple points.
Amongst these codes, we observed that eight of them are directly or indirectly related to heart attack. 411.89 is directly related, as ischemic heart disease [feske2021ischemic] often leads to myocardial infarction. 429.9 is also directly linked, as it encompasses various heart conditions, including those leading to heart attacks. Indirectly related codes include 428.20 since heart failure can result from repeated heart attacks or prolonged ischemia; 398.90, as valve damage from rheumatic fever increases the risk of ischemic events [remenyi2016valvular]; 440.21 and 440.4 since atherosclerosis is a major cause of coronary artery disease and heart attacks [falk2006pathogenesis]; 429.83, Takotsubo syndrome, which mimics a heart attack and is triggered by extreme stress [butt2022long]; and 996.61, as infections from cardiac implants can lead to complications affecting heart function [klug2007risk]. These conditions either directly cause myocardial infarction or indicate underlying cardiovascular issues that increase heart attack risk.
@@
TABLE V: ICD-9 Codes and Their Descriptions for the teal points.
Similarly, the cluster of points colored in teal primarily contains ICD-9 codes associated with mental health and neurological issues. A detailed description of the ICD-9 codes and their meanings is given in Table V. 386.11 is directly related, as benign paroxysmal positional vertigo (BPPV) affects the vestibular system, leading to dizziness and balance issues [bhattacharyya2008clinical]. 357.0 and 357.4 involve polyneuropathies [sommer2018polyneuropathies], which affect the peripheral nervous system and can result in weakness, numbness, or pain. 784.69 is linked to speech disturbances, often indicative of neurological disorders. Mental health conditions are represented by 296.44 and 296.42, which describe different severities of bipolar I disorder, impacting mood and cognition. 295.40 covers schizophreniform disorder, a condition with symptoms resembling schizophrenia. Substance-related disorders, such as 305.43 (opioid abuse with other drug abuse) and 305.30 (cannabis abuse), highlight the intersection between mental health and addiction. 437.8, which includes cerebrovascular diseases, may contribute to cognitive impairment and neurological deficits. These conditions directly or indirectly affect brain function, mental well-being, and the nervous system.
@@ VI Conclusion
VII Acknowledgements
This project is partially supported by the NSF Career Award IIS 2442159. Its contents are solely the responsibility of the authors and do not necessarily represent the official views of the NSF.